# Learning thin deformable object manipulation with a multi-sensory integrated soft hand

### v1

@@ IV-F Visualization the role of multisensory information
In the t-SNE visualization for F/T sensing (Fig. 8A), the close and overlapping embeddings suggest that different objects share similar properties in the F/T feature space. Fig. 8(D to E) reveal a clear correlation between F/T sensing and the number of layers in the book and fabric booklet, indicating that F/T sensing primarily reflects physical interaction properties such as pressure, weight, or finger positioning, rather than the material properties.
@@ IV-G Discussion
We see several limitations and opportunities for future work. First, our learned policy relies on multisensory information obtained before manipulation. This is a significant advantage, enabling one-shot manipulation and enhanced efficiency. However, the reliance on pre-obtained multisensory information has its inherent limitations. A major opportunity for future research would use the multisensory information gathered after manipulation (e.g. after the two fingers are closed). This could allow for active determination of whether multiple layers have been grasped and guide subsequent actions of the robot. Another limitation is that when one layer has significantly lower friction than the other in stacked objects, one-handed manipulation is not effective enough for singulation. Even humans will encounter difficulties in such conditions. Just as humans would use both hands to coordinate manipulation, one opportunity for future work is to learn a dual-hand system to address this situation.
Furthermore, our current manipulation policy obtains visual perception from the depth image without fully utilizing the color image. The learned policy is independent of information such as object color, which eliminates the need to retrain or fine-tune across different environments. However, depth information ignores details present in color images that may offer insights into materials and textures. Integrating the RGB image into the network potentially improves policy performance. Moreover, our slip module is simple and effective in predicting slip points and is learned prior to RL policy training. However, the slip module may require additional training data when faced with novel objects of different structures or shapes. This limitation stems from its inherent lack of world knowledge and inability to engage in physical reasoning. For instance, it cannot understand why a slip point would be more effective on the upper-right corner of a book than on the spine. Such knowledge usually reflects humans’ high-level cognitive intelligence and is distinct from the focus of this paper, which is on low-level sensory-motor intelligence. Looking ahead, an interesting direction for future research is to leverage large foundation models [60, 61] to learn high-level planning and reasoning about where to perform the slip motion. Such models have built extensive world knowledge and representations, enabling direct generalization to a wider range of objects and facilitating robot deployment in more complex environments. Additionally, the in-hand manipulation model can be learned to align the object to the appropriate position after grasping, which could be widely used in the subsequent folding of clothes or origami tasks.
@@ V Conclusion
This study presents the method for learning to singulate and grasp thin, deformable objects. The learned policy drives our robot equipped with soft pneumatic grippers, achieving a significant leap over past work in terms of graspable object variety and performance. Notably, our training uses only printer paper and winter fabric on flat surfaces. However, the robot successfully handles various deformable objects with different physical properties (coated paper, hotel towels, T-shirts), hybrid materials (fabrics separated by introduction pages or spring pancakes separated by baking paper), and gravitational effects. Our results suggest that the extraordinary complexity of manipulating thin, deformable objects can be tamed with embodied multisensory integration and compliance in the soft gripper. By enabling the robot to adapt to the complex dynamics of deformable objects without relying on precise control, our approach opens new possibilities for robots to operate in unstructured environments and perform tasks that require delicate handling. Future work may involve extending this framework to more complex manipulation tasks and further improving the efficiency and generalization of the learning process.
VI Acknowledgments
Funding: the research is supported by the HKSAR (project: ITS/036/21FP) and the Foshan HKUST (project: FSUST21-HKUST03B, FSUST21- FYTRI04B). Data availability: Videos of experiments, the hardware models, instruction, data, and code will be posted at https://robotll.github.io/LTDOM/.
VII Appendix
@@
Sections S1 to S9
Figs. S1 to S6
Tables S1 to S4
Other Supplementary materials, including Movie 1, Movies S1 to S2 are at https://robotll.github.io/LTDOM/
@@
Quantitative results from both sets of experiments are presented in Tab. S1. Notably, the above test settings significantly extend beyond the scope of existing literature on deformable object manipulation, which usually assumes that only a piece of clothing or a bag is on the table. The outcomes of these exploratory trials underscore the robustness and adaptability of our methodology, demonstrating its promise for real-world applications in clothing and bag manipulations, even under complex conditions.
@@
Fig. S2: Analysis of failure cases and touch sensor design. (A) A failure case where the friction between layer B and layer C is so low that both layer A and layer B lift. (B) A failure case where layer A adheres to layer B, resulting in the unintended lifting of both layers. (C) The design of our touch sensor, which includes the touch, imaging, and support modules. (D) This image shows the robotic gripper equipped with the touch sensor, highlighting the sensing area. (E) Raw image from the sensor. (F) Stitched image created from 15 tiles of the raw image. (G) The final obtained deformation field.
@@
The second cause of failure, as illustrated in Fig. S2B, arises when layers are stuck together. This is commonly observed in contexts involving paper where even manual separation requires considerable effort. The adhesion often results from microscopic structural factors or external variables like ink, printing, and Van der Waals forces. Under such conditions, the frictional force between layers A and B is significantly greater than between layers B and C, leading to multiple layers being grasped together. As discussed in the experiments of book page-turning with printer paper, we attempted 200 flips and succeeded 188 times, yielding a success rate of 94%. Among the 12 unsuccessful attempts, 11 were attributed to tight adhesion between pages and one to noise errors in the depth camera reading.
@@
We used a vision-based tactile sensor of fingertip dimensions based on [54]. While sensors like GelSight [49] and DIGIT [50] can offer high-resolution tactile feedback, they are constrained by their larger dimensions and flat geometry; hence they are hard to integrate into a space the size of a human fingertip. In contrast, our sensor is able to provide a large field of view in a compact form factor, benefiting from the compound-eye imaging system [54], making it ideal for integration into the minuscule fingertips of soft grippers.
Our tactile sensor consists of three crucial components, as illustrated in Fig. S2C: the touch, imaging, and support modules. The touch module features a curved elastomer layer designed to conform to the fingertip of our soft gripper, and an acrylic plate. The elastomer layer is constructed by a transparent layer embedded with a random color pattern, a white protective layer, and a black layer designed to stabilize light intensity variations. The imaging sub-system beneath the curved touching layers includes a microlens array, a pinhole structure to define the aperture, and a Complementary Metal Oxide Semiconductor (CMOS) image sensor. Enclosing these is the support module, a 3D-printed shell that provides the structural integrity required for the effective functioning of the tactile sensor. Fig. S2D displays the sensor whens mounted on the soft finger. We obtain the touch information with the following steps. First, we crop 15 tiles from the raw image (Fig. S2E), which are subsequently stitched together (Fig. S2F). Next, we employ the same algorithm in [54] to process the combined image and obtain the deformation field (Fig. S2G).
@@
The observation vectors are defined in Tab. S3. Visual observation $o_{t}^{vis}$ is defined as a $40\times 40$ depth image cropped from the region surrounding the slip point. Tactile information $o_{t}^{ind},o_{t}^{thu}$ at each fingertip is encapsulated in a $25\times 25\times 3$ matrix, which represents the deformation field in the $\alpha$, $\beta$, and $\gamma$ directions, and samples 625 points across the tactile area.
In the fine-grained action space of the inner loop, $z_{t}$ ranges from -3 mm to 3 mm at 1 mm intervals, approximating the thickness of 60 sheets of printer paper or 10 layers of winter fabric. $x_{t}$ varies between -7.5 mm and 7.5 mm, with a 5 mm interval, and $\theta_{t}$ is confined between 0° and 3°, at 1° intervals. Conversely, the ranges in the coarse action space are 2.5 times larger than those in the fine action space, with the same discrete number. Therefore, $z_{t}$ extends from -7.5 mm to 7.5 mm, $x_{t}$ spans from -18.7 mm to 18.7 mm, and $\theta_{t}$ ranges between 0° and 7.5°.
Fig. S3: Annotated images and models of robot hand. (A to C) Annotation on three kinds of objects used for learning slip network. (D) 3D model for the two-fingered soft robot hand. (E) Internal air chamber structure of the finger.
@@
The network within the slip module is designed as a fully convolutional neural network (FCN). Specifically, we discretize the image rotation within a range of $(-90^{\circ},90^{\circ})$ at intervals of $15^{\circ}$, corresponding to slip directions in the environment. To ensure the object remains within the frame post-rotation, we pad the image prior to adjusting its resolution to $640\times 480$. Then the slip network utilizes a ResNet-18 backbone [62] for feature extraction. Upon extracting the features, the network incorporates three $1\times 1$ convolution layers. The initial layer reduces the 512 channels from ResNet-18 outputs to 128 channels, followed by 2x bilinear upsampling. The next layer further reduces the channels to 32, followed by another 2x bilinear upsampling. The last layer maps these 32 channels to a $640\times 480$ probability map. The slip network is optimized using the Adam optimizer [59], with a learning rate of 0.001 and a binary cross entropy loss function. The training dataset is extended by data augmentation from manually annotated images (Fig. S3 A to C).
@@
TABLE S3: Dimensionality of observations.
{tblr}
visual observation $o_{t}^{vis}$ 40x40
tactile observation $o_{t}^{ind}$ 25x25x3
$o_{t}^{thu}$ 25x25x3
muscle proprioception $o_{t}^{pro}$ 6
auxiliary information $o_{t}^{aux}$ 2
@@
For base encoders, the visual encoder $g_{\text{vis}}$ is designed as a Convolutional Neural Network (CNN), consisting of two convolutional layers with 16 and 32 filters, each employing a 7x7 kernel and a stride of 2. The layers are followed by Rectified Linear Unit (ReLU) [63] activation functions and an adaptive average pooling layer. The touch encoder $g_{\text{ind}}$ and $g_{\text{thu}}$ are also modeled as CNNs, mirroring the visual encoder in kernel size, stride, and padding. The muscle proprioception encoder $g_{\text{prop}}$ and auxiliary state encoder $g_{\text{aux}}$ are constructed as linear layers containing 32 hidden units. In the initial phase, each observation element is transformed into a 32-dimensional latent vector using base encoders, which are then fed into the multisensory encoder.
@@
Finally, the 64-dimensional output of the transformer encoder is fed into the MLP. The final MLP block has two hidden layers with $\{64,1\}$ hidden units for the outer loop and $\{64,4\}$ hidden units for the inner loop. The details of the hyper-parameters for policy training can be found in Tab. S2.
Fig. S4: Comparison of ablation model architectures. (A) Our method architecture incorporates inputs from various sensor modalities. (B) OnlyVision architecture relies solely on visual input. (C) NoTouch architecture excludes touch input but includes other modalities. (D) SingleLoop architecture is a version without inner-outer loop learning process; thus, it does not incorporate auxiliary status inputs. (E) NoAttention architecture employs an MLP without the attention mechanism for processing sensor inputs.
@@
  • SingleLoop (SL): This method only uses the inner loop to learn manipulation policy in the entire action space, eliminating the need for auxiliary observation input, as depicted in Fig. S4D. Specifically, the range for $z_{t}$ is from -7.5 mm to 7.5 mm, with an interval of 1 mm. $x_{t}$ varies between -20 mm and 20 mm, with a step size of 5 mm, while $\theta_{t}$ is confined between $0\degree$ and $7.5\degree$, spaced at $1.5\degree$ intervals. This ablation model serves to assess the efficiency of the proposed dual-loop learning structure under real-world training conditions;
@@
  • NoAttention (NA): We concatenate the latent feature vectors from the basic encoder and replace the multisensory encoder with an MLP block (Fig. S4E). This setup aims to evaluate the effectiveness of the proposed attention-based multisensory encoder when processing multisensory information.
Fig. S5: Experiments of including different gravitational effects in policy training. (A and B) Success rate offsets, calculated as the difference in performance of test objects between policies trained on various workspace tilt angles and the success rate of our original policy, which was trained exclusively on a horizontal workspace. (C) Training curves for policy learned with different tilt angles of the workspace. The time when humans exchange objects is not included. Training stops when performance no longer increases.
S9 Experiments of training policies on different gravitational conditions
In this section, we explore the policy performance trained under conditions that include different gravity effects. We compared three cases: our original policy trained only on the horizontal plane (denoted as OnlyZ) and a policy trained at workspace tilt angles including 0 and 30 degrees (denoted as Zand30) and 0 and 60 degrees (denoted as Zand60). The training objects for all cases are the same, including printer books and winter suit fabrics. Unseen objects for testing are those cases in Fig. 3D, excluding the training objects. During training, the two robotic arms are each responsible for a tilt angle. After every five attempts, the researcher exchanges objects.
Fig. S5A and 5B show the difference in performance when training includes different gravity effects, compared to OnlyZ. The results show that Zand60 significantly improves the average performance in the test under the 30 and 60-degree conditions (Fig. S5B). This proves that including different gravity effects in the training can enhance the overall performance. Meanwhile, Zand30 also achieved performance improvement under corresponding conditions, although the improvement was smaller (Fig. S5A). This reflects that objects on the physical characteristics of a workspace with a 30-degree tilt angle are closer to those of a 0-degree workspace.
Fig. S5C shows the training curves of these three cases. Including different workspace tilt angles increases the diversity of situations encountered during the training phase. This variation demands more exploration from the RL agent, thus extending the training time to reach convergence.
These results also demonstrate the potential that, when hardware conditions permit, our approach can be extended to more parallel environments. This can not only significantly accelerate the learning process but also cover a wider range of training objects and gravitational effects, further improving the robot’s performance.

### v2

@@ IV-F Visualization the role of multisensory information
In the t-SNE visualization for F/T sensing (Fig. 8A), the close and overlapping embeddings suggest that different objects share similar properties in the F/T feature space. Fig. 8(D to E) reveals a clear correlation between F/T sensing and the number of layers in the book and fabric booklet, indicating that F/T sensing primarily reflects physical interaction properties such as pressure, weight, or finger positioning, rather than the material properties.
@@ IV-G Discussion
We see several limitations and opportunities for future work. First, our learned policy relies on multisensory information obtained before manipulation. This is a significant advantage, enabling one-shot manipulation and enhanced efficiency. However, the reliance on pre-obtained multisensory information has its inherent limitations. A major opportunity for future research would use the multisensory information gathered after manipulation (e.g., after the two fingers are closed). This could allow for active determination of whether multiple layers have been grasped and guide subsequent actions of the robot. Another limitation is that when one layer has significantly lower friction than the other in stacked objects, one-handed manipulation is not effective enough for singulation. Even humans will encounter difficulties in such conditions. Just as humans would use both hands to coordinate manipulation, one opportunity for future work is to learn a dual-hand system to address this situation.
Furthermore, our current manipulation policy obtains visual perception from the depth image without fully utilizing the color image. The learned policy is independent of information such as object color, which eliminates the need to retrain or fine-tune across different environments. However, depth information ignores details present in color images that may offer insights into materials and textures. Integrating the RGB image into the network potentially improves policy performance. Moreover, our slip module is simple and effective in predicting slip points and is learned prior to RL policy training. However, the slip module may require additional training data when faced with novel objects of different structures or shapes. This limitation stems from its inherent lack of world knowledge and inability to engage in physical reasoning. For instance, it cannot understand why a slip point would be more effective on the upper-right corner of a book than on the spine. Such knowledge usually reflects humans’ high-level cognitive intelligence and is distinct from the focus of this paper, which is on low-level sensory-motor intelligence. Looking ahead, an interesting direction for future research is to leverage large foundation models [60, 61] to learn high-level planning and reasoning about where to perform the slip motion. Such models have built extensive world knowledge and representations, enabling direct generalization to a wider range of objects and facilitating robot deployment in more complex environments. Additionally, our current approach demonstrates the power of passive compliance to reduce the reliance on precise perception and control. In future work, combining the passive compliance (for safety and adaptability) with impedance-like active control (for precision when needed) could further improve performance.
@@ V Conclusion
This study presents the method for learning to singulate and grasp thin, deformable objects. The learned policy drives our robot equipped with soft pneumatic grippers, achieving a significant leap over past work in terms of graspable object variety and performance. Notably, our training uses only printer paper and winter fabric on flat surfaces. However, the robot successfully handles various deformable objects with different physical properties (coated paper, hotel towels, T-shirts), hybrid materials (fabrics separated by pages or spring pancakes separated by baking paper), and gravitational effects. Our results suggest that the extraordinary complexity of manipulating thin, deformable objects can be tamed with embodied multisensory integration and compliance in the soft gripper. By enabling the robot to adapt to the complex dynamics of deformable objects without relying on precise control, our approach opens new possibilities for robots to operate in unstructured environments and perform tasks that require delicate handling. Future work may involve extending this framework to more complex manipulation tasks and further improving the efficiency and generalization of the learning process.
VI Appendix
@@
Sections S1 to S10
Figs. S1 to S4
Tables S1 to S7
Other Supplementary materials, including Movie 1, Movies S1 to S2, part of the hardware models, instructions, data, and code, will be at https://robotll.github.io/LTDOM/
@@
Quantitative results from both sets of experiments are presented in Table S1. Notably, the above test settings significantly extend beyond the scope of existing literature on deformable object manipulation, which usually assumes that only a piece of clothing or a bag is on the table. The outcomes of these exploratory trials underscore the robustness and adaptability of our methodology, demonstrating its promise for real-world applications in clothing and bag manipulations, even under complex conditions.
@@
Fig. S2: Analysis of failure cases and touch sensor design. (A) A failure case where the friction between layer B and layer C is so low that both layer A and layer B lift. (B) A failure case where layer A adheres to layer B, resulting in the unintended lifting of both layers. (C) The design of our touch sensor, which includes the touch, imaging, and support modules. (D) This image shows the robotic gripper equipped with the touch sensor, highlighting the sensing area. (E) Raw image from the sensor. (F) Stitched image created from 15 tiles of the raw image. (G) The final deformation field.
@@
The second cause of failure, as illustrated in Fig. S2B, arises when layers are stuck together. This is commonly observed in contexts involving paper, where even manual separation requires considerable effort. The adhesion often results from microscopic structural factors or external variables like ink, printing, and Van der Waals forces. Under such conditions, the frictional force between layers A and B is significantly greater than between layers B and C, leading to multiple layers being grasped together. As discussed in the experiments of the book page turning with printer paper, we attempted 200 flips and succeeded 188 times, yielding a success rate of 94%. Among the 12 unsuccessful attempts, 11 were attributed to tight adhesion between pages and one to noise errors in the depth camera reading.
@@
We used a vision-based tactile sensor of fingertip dimensions based on [54]. While sensors like GelSight [49] and DIGIT [50] can offer high-resolution tactile feedback, they are constrained by their larger dimensions and flat geometry; hence, they are hard to integrate into a space the size of a human fingertip. In contrast, our sensor is able to provide a large field of view in a compact form factor, benefiting from the compound-eye imaging system [54], making it ideal for integration into the minuscule fingertips of soft grippers.
Our tactile sensor consists of three crucial components, as illustrated in Fig. S2C: the touch, imaging, and support modules. The touch module features a curved elastomer layer designed to conform to the fingertip of our soft gripper, and an acrylic plate. The elastomer layer is constructed by a transparent layer embedded with a random color pattern, a white protective layer, and a black layer designed to stabilize light intensity variations. The imaging sub-system beneath the curved touching layers includes a microlens array, a pinhole structure to define the aperture, and a Complementary Metal Oxide Semiconductor (CMOS) image sensor. Enclosing these is the support module, a 3D-printed shell that provides the structural integrity required for the effective functioning of the tactile sensor. Fig. S2D displays the sensor when mounted on the soft finger. We obtain the touch information with the following steps. First, we crop 15 tiles from the raw image (Fig. S2E), which are subsequently stitched together (Fig. S2F). Next, we employ the same algorithm in [54] to process the combined image and obtain the deformation field (Fig. S2G).
@@
The observation vectors are defined in Table S2. Visual observation $o_{t}^{vis}$ is defined as a $40\times 40$ depth image cropped from the region surrounding the slip point. Tactile information $o_{t}^{ind},o_{t}^{thu}$ at each fingertip is encapsulated in a $25\times 25\times 3$ matrix, which represents the deformation field in the $\alpha$, $\beta$, and $\gamma$ directions, and samples 625 points across the tactile area.
In the fine-grained action space of the inner loop, $z_{t}$ ranges from -3 mm to 3 mm at 1 mm intervals, approximating the thickness of 60 sheets of printer paper or 10 layers of winter fabric. $x_{t}$ varies between -7.5 mm and 7.5 mm, with a 5 mm interval, and $\theta_{t}$ is confined between 0° and 3°, at 1° intervals. Conversely, the ranges in the coarse action space are 2.5 times larger than those in the fine action space, with the same discrete number. Therefore, $z_{t}$ extends from -7.5 mm to 7.5 mm, with a 1 mm interval, $x_{t}$ spans from -18.7 mm to 18.7 mm, with a 5 mm interval, and $\theta_{t}$ ranges between 0° and 7.5°, at 1° interval.
Fig. S3: Annotated images and models of robot hand. (A to C) Annotation on three kinds of objects used for learning the slip network. (D) 3D model for the two-fingered soft robot hand. (E) Internal air chamber structure of the finger.
@@
The network within the slip module is designed as a fully convolutional neural network (FCN). Specifically, we discretize the image rotation within a range of $(-90^{\circ},90^{\circ})$ at intervals of $15^{\circ}$, corresponding to slip directions in the environment. To ensure the object remains within the frame post-rotation, we pad the image prior to adjusting its resolution to $640\times 480\times 1$. Then the slip network utilizes a ResNet-18 backbone [62] for feature extraction. Upon extracting the features, the network incorporates three $1\times 1$ convolution layers. The initial layer reduces the 512 channels from ResNet-18 outputs to 128 channels, followed by 2x bilinear upsampling. The next layer further reduces the channels to 32, followed by another 2x bilinear upsampling. The last layer maps these 32 channels to a $640\times 480$ probability map. The slip network is optimized using the Adam optimizer [59], with a learning rate of 0.001 and a binary cross-entropy loss function. The training dataset is extended by data augmentation from manually annotated images (Fig. S3 A to C).
TABLE S2: Dimensionality of observations.
{tblr}
visual observation $o_{t}^{vis}$ 40x40
tactile observation $o_{t}^{ind}$ 25x25x3
$o_{t}^{thu}$ 25x25x3
muscle proprioception $o_{t}^{pro}$ 6
auxiliary information $o_{t}^{aux}$ 2
@@
For base encoders, the visual encoder $g_{\text{vis}}$ is designed as a Convolutional Neural Network (CNN), consisting of two convolutional layers with 16 and 32 filters, each employing a 7x7 kernel and a stride of 2. The layers are followed by Rectified Linear Unit (ReLU) [63] activation functions and an adaptive average pooling layer. The touch encoder $g_{\text{ind}}$ and $g_{\text{thu}}$ are also modeled as CNNs, mirroring the visual encoder in kernel size, stride, and padding. The muscle proprioception encoder $g_{\text{pro}}$ and auxiliary state encoder $g_{\text{aux}}$ are constructed as linear layers containing 32 hidden units. In the initial phase, each observation element is transformed into a 32-dimensional latent vector using base encoders, which are then fed into the multisensory encoder.
@@
Finally, the 64-dimensional output of the transformer encoder is fed into the MLP. The final MLP block has two hidden layers with $\{64,1\}$ hidden units for the outer loop and $\{64,4\}$ hidden units for the inner loop. The details of the hyperparameters for policy training can be found in Table S3.
Fig. S4: Comparison of ablation model architectures. (A) Our method architecture incorporates inputs from various sensor modalities. (B) OnlyVision architecture relies solely on visual input. (C) NoTouch architecture excludes touch input but includes other modalities. (D) NoForce architecture excludes the F/T sensor input but includes other modalities. (E) SingleLoop architecture is a version without the inner-outer loop learning process; thus, it does not incorporate auxiliary status inputs. (F) NoAttention architecture employs an MLP without the attention mechanism for processing sensor inputs.
@@
  • NoForce (NF): The observation is exclusively based on visual and tactile perception to validate the importance of force and torque information. Refer to Fig. S4D for a visual representation of this ablation model architecture;
@@
  • SingleLoop (SL): This method only uses the inner loop to learn manipulation policy in the entire action space, eliminating the need for auxiliary observation input, as depicted in Fig. S4E. Specifically, the range for $z_{t}$ is from -7.5 mm to 7.5 mm, with an interval of 1 mm. $x_{t}$ varies between -18.7 mm and 18.7 mm, with a step size of 5 mm, while $\theta_{t}$ is confined between $0\degree$ and $7.5\degree$, spaced at $1\degree$ intervals. This ablation model serves to assess the efficiency of the proposed dual-loop learning structure under real-world training conditions;
@@
  • NoAttention (NA): We concatenate the latent feature vectors from the basic encoder and replace the multisensory encoder with an MLP block (Fig. S4F). This setup aims to evaluate the effectiveness of the proposed attention-based multisensory encoder when processing multisensory information.
TABLE S4: Comparison of our gripper with prior designs for thin-object manipulation and general soft robotic hands.
{tblr}
cells = c, vline2 = 2-7, hline1-2,8 = -, Feature & Ours Flipbot[9] RoTipBot[10] Tirumala et al.[8] Zhang et al.[41] Truby et al.[42] Puhlmann et al.[43]
Gripper type Soft Soft Rigid w/ soft tactile Rigid Soft Soft Soft
Actuation Pneumatic Pneumatic Motorized Motorized Pneumatic Pneumatic Pneumatic
Compliance High High Low Low High High High
Control complexity Low Low Medium Low High Medium High
Integrated sensors Fingertip tactile, F/T F/T Fingertip tactile Fingertip tactile Palm tactile    Ionogel -
Soft structure fabrication Easy Easy - - Complex Complex Medium
S9 Novelty and Rationale of the Gripper Design
Our gripper design is based on the well-established fast-PneuNet [53] architecture but uniquely integrates several key features tailored explicitly for manipulating thin, deformable objects. These key features include underactuation, passive compliance, and multimodal sensing via fingertip tactile sensors combined with a wrist-mounted force/torque sensor.
Compared to other grippers used in literature for grasping thin, deformable objects, while [10, 8] utilizes fingertip tactile sensors as ours, its rigid gripper lacks the whole-finger passive compliance intrinsic to our soft fingers. We have empirically demonstrated in our experiments (Section IV-D) the substantial contribution of passive compliance in the successful manipulation of thin, deformable objects.
Furthermore, in comparison to other soft robotic hands, our design emphasizes simplicity and practical usability. Each finger utilizes only a single pneumatic channel for bending. This underactuation significantly reduces the control complexity, making it far tractable than controlling multiple joints in a rigid or more complex multi-chambered soft hand.
For example, [41] focuses on achieving human-like dexterity through coordinated palm-finger movements and high-density tactile sensing in the palm. Although highly capable, such control complexity is not strictly necessary for the singulation and grasping of thin, deformable objects. Similarly, [42] relies on discrete, complex actuation modes that significantly increase fabrication and control overhead. Given our model-free RL framework, which learns directly in the real world, minimizing control complexity is crucial for efficient training and generalization. Our hand strikes a balance between sensory richness and control simplicity, enabling effective performance without over-engineering.
Another critical advantage of our gripper is its manufacturability and accessibility. The soft fingers can be fabricated within approximately 12 hours using consumer-grade 3D printers (e.g., Bambu P1P, priced approximately \$500-\$700) and widely available thermoplastic polyurethane (TPU) filament (approximately \$40). This straightforward manufacturing approach eliminates the need for specialized processes like silicone casting [41] or embedded 3D printing [42], significantly reducing barriers for adoption within the robotics research community. While we chose to integrate our own fingertip tactile sensors for convenience, it is worth noting that the geometry of the finger is also easily modifiable due to its reproducibility and ease of fabrication, allowing for easy adaptation to a variety of commercially available or custom sensors as needed.
In summary, the novelty of our gripper design lies in its specific combination of underactuation, passive compliance, and multimodal sensory integration, all tailored for the challenges of manipulating thin deformable objects using reinforcement learning. This unique integration reduces control complexity, enhances robustness and adaptability, provides rich sensory feedback essential for our slip module, and supports the ”imprecise dexterity” required for manipulating thin, deformable objects. Detailed comparisons across different robotic hands are summarized in Table S4.
TABLE S5: Success rates (%) across different object sizes. The medium size corresponds to the objects utilized in the experiments presented in the main text.
{tblr}
columneven = c, column3 = c, hline1-2,5 = -, Object Type & Small Medium Large
Printer Paper (A6, A4, A2) 90 91 89
Towel (15x15, 30x30, 45x45 cm) 98 98 97
T-shirt (XXS, S, L) 84 86 87
@@
S10 Discussion of action space granularity generalization
A pertinent question regarding the dual-loop learning framework presented in this work is whether its chosen coarse and fine action space granularities are effective only for the particular type and size of thin deformable objects tested in the experiments, or if they have broader applicability.
The defined action space granularities can generalize across various object sizes due to the inherently localized nature of interactions involved in manipulating thin, deformable objects. Analogous to human manipulation tasks, such as page-turning, interactions primarily occur at specific localized regions (e.g., corners or edges) regardless of the overall object dimensions. Hence, substantial variations in object sizes minimally affect the fundamental localized deformation and interaction dynamics, allowing the established coarse and fine action granularities to remain effective.
To quantitatively validate this assertion, we performed supplementary experiments involving objects of diverse dimensions, including printer paper sheets (A2, A4, A6), hotel towels (15 cm × 15 cm, 30 cm × 30 cm, 45 cm × 45 cm), and T-shirts (XXS, S, L). Results summarized in Table S5 demonstrate that size variations minimally influence the effectiveness of the chosen action granularities. For example, the page-turning task consistently achieved high success rates (approximately 90%) across all paper sizes. Comparable success rates were observed in grasping tasks involving towels and T-shirts, confirming robust generalization across a broad spectrum of object sizes.
In principle, the current action granularity should remain effective for any unseen object whose physical properties fall within the range of the materials we tested (e.g., similar surface friction or bending stiffness). Our policy does not rely on predefined strategies for specific materials but instead learns to associate different sensory signals with the required actions (or action granularities).
The learned policy used for testing is trained solely on printer paper and winter suit fabric, all without any modification to granularity. As shown in our results, the granularity settings have exhibited strong generalization across various unseen object types, including coated paper, plastic paper, summer suit fabrics, hotel towels, T-shirts, and hybrid paper/fabric stacks. These objects span a significant range of thicknesses, stiffnesses, surface textures, and friction coefficients, representing typical thin deformable objects encountered daily. This empirical evidence supports the hypothesis that, for unseen objects with similar physical properties to those we tested, our action space granularities will remain effective.